OpenStudy (anonymous):

The average of four students' scores on a test was 70. The sum of seven other students' scores on the same math test was 160. What was the average score for these eleven students together?

OpenStudy (anonymous):

The answer is 40, but for my work i just did 160+70=230 230/2=115

OpenStudy (anonymous):

well the average of four students' scores is 70, so the total for the four students is 70 x 4 = 280 add this to the total of the other seven students 280 + 160 = 440 divide that by the eleven students and you can find the average. 440/11 = answer

OpenStudy (anonymous):

The answer you are looking for is: \[Average = \frac{4 \times 70 + 160 }{11}\] Solve this...

OpenStudy (anonymous):

thanks to both of you. :)

OpenStudy (anonymous):

Welcome..